OpenStudy (anonymous):

1. v=v0+at 2.x=x0+v0t+(1/2)at^2 For problem solving it is useful to eliminate a by substituting for a from 1 Into 2. x=x0+(1/2)(v0+v)t For problem solving eliminate t in 2 by substituting for t from 1. v^2=v0^2+2a(x-x0) I keep messing this up. For the first one I keep geting x=x0+(1/2)vt

OpenStudy (lgbasallote):

1. v = v0 + at v - v0 = at v - v 0 = a ------ t agree?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so that would give you x = x0 + v0t + (v - v0)t^2 / 2t so you have \[\implies x =x_0 + v_0t + \frac{(v-v_0)t}2\]

OpenStudy (lgbasallote):

then if you factor out t.. \[\implies x = x_0 + t\;(v_0 + \frac{v - v_0}2)\] convert to similar denominators.. \[\implies x = x_0 + t\; (\frac{2V_0}2 - \frac{v-v_0}2)\] following?

OpenStudy (anonymous):

yeah I'm following.

OpenStudy (lgbasallote):

typo... that should be \[\implies x = x_0 + t \; (\frac{2V_0}2 + \frac{v-v_0}2)\]

OpenStudy (lgbasallote):

so if you combine them you get \[\implies x = x_0 + t \; (\frac{v + v_0}2)\] that was your question right?

OpenStudy (anonymous):

For the second one, would it be easier to use the answer for the first one to find. v^2=v0^2+2a(x-x0)

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

The second part is: For problem solving eliminate t in 2 by substituting for t from 1. v^2=v0^2+2a(x-x0)

OpenStudy (anonymous):

Just use x=x0+(1/2)(v0+v)t and t=(v-v0)/a

OpenStudy (lgbasallote):

sorry i have to go now. important business call

OpenStudy (lgbasallote):

but you should solve for t in v = v_0t + at then substitute that t into x = x_0 + v_0 t + 1/2 at^2 that should give you the answer